.

TimestampEvent TimeAnnotationPersonUTC TimestampLast modified

.

12/4/2012 15:39:00today at 3Tested this scriptRyan Hunt12/4/2012 21:39:002013-01-23

.

12/4/2012 15:46:06Loaded the cell with Helium
Ryan Hunt12/4/2012 21:46:06

.

12/4/2012 15:15:00Manually entered as a test
Ryan Hunt12/4/2012 21:15:002012-12-04

.

12/4/2012 15:51:2015:15:00Malachi cooled, vacuumed the cell, filled with 6.4 Bar He, and then turned the power up to 48 W so we can do a pressure drop experiment in Helium.
Ryan Hunt12/4/2012 21:51:20

.

12/4/2012 16:48:01Released helium from cell down to ~7 bar.Malachi Heder12/4/2012 22:48:01

.

12/4/2012 18:03:26Released helium to ~ 6 barMalachi Heder 12/5/2012 0:03:26

.

12/4/2012 19:30:06Released helium down to ~5 barMalachi Heder12/5/2012 1:30:06

.

12/4/2012 20:56:14Released pressure to ~4 barMalachi Heder12/5/2012 2:56:14

.

12/4/2012 22:46:01Released helium pressure to ~3 barMalachi Heder12/5/2012 4:46:01

.

12/5/2012 8:46:38Released helium pressure to ~2 barMalachi Heder12/5/2012 14:46:38

.

12/5/2012 9:22:19Released helium pressure to ~1 bar.Malachi Heder12/5/2012 15:22:19

.

12/7/2012 9:04:49Updating firmware, then setting power to 48 wattsMalachi Heder12/7/2012 15:04:49

.

12/7/2012 12:54:57Bringing pressure to .5 bar.Malachi Heder12/7/2012 18:54:57

.

12/7/2012 12:59:42Over shot mark, vacuum and trying again.Malachi Heder12/7/2012 18:59:42

.

12/7/2012 13:16:1612/6/2012 17:00:00I loading new firmware turned the voltage control off. Malachi Heder12/7/2012 19:16:16

.

12/10/2012 9:40:39Turned cell off, preparing to swap cells.Malachi Heder12/10/2012 15:40:39

.

12/10/2012 9:55:38Stopped data collection to swap test cells.Malachi Heder12/10/2012 15:55:38

.

12/10/2012 10:38:00Data collection running, checking all sensors.Malachi Heder12/10/2012 16:38:00

.

12/10/2012 10:46:42Calibrating pressure relief valve to 8 bar.Malachi Heder12/10/2012 16:46:42

.

12/10/2012 11:33:53Checking vacuum.Malachi Heder12/10/2012 17:33:53

.

12/10/2012 15:26:38Filling with 6.4 bar hydrogen to test for leaks and prep for first pressure sweep.Malachi Heder12/10/2012 21:26:38

.

12/11/2012 9:48:21Swing down the pressure to 6 bar. Continuing pressure sweep at 48 watts.Malachi Heder12/11/2012 15:48:21

.

12/11/2012 10:37:00Increased power to 48 watts.Malachi Heder12/11/2012 16:37:00

.

12/11/2012 11:18:33Reduced pressure to 5 bar, increased power to 49 watts manually.Malachi Heder12/11/2012 17:18:33

.

12/11/2012 13:51:56Just decreased Hydrogen pressure to 3.4 (tried for 4 bar, but the new needle valve on this cell is tricky)Ryan12/11/2012 19:51:56

.

12/11/2012 15:09:07Released pressure to 3 bar.Malachi Heder 12/11/2012 21:09:07

.

12/11/2012 15:59:52Reduced pressure to 1.9 bar for almost last stepRyan12/11/2012 21:59:52

.

12/12/2012 9:39:19Decreased H2 pressure to just under 1 bar Ryan Hunt12/12/2012 15:39:19

.

12/12/2012 13:00:06Turned off power. Malachi Heder12/12/2012 19:00:06

.

12/14/2012 16:05:38Vacuuming cell, new additions include mineral wool.Malachi Heder12/14/2012 22:05:38

.

12/14/2012 16:26:11Putting positive pressure of hydrogen in cell to check for leaks. Malachi Heder12/14/2012 22:26:11

.

12/14/2012 16:30:08No leaks detected on cell, continuing to pressurize to 8 bar for pressure sweep test.Malachi Heder12/14/2012 22:30:08

.

12/14/2012 16:31:23Check that, pressurizing to 6.4 bar to allow for thermal expansion of gas.Malachi Heder12/14/2012 22:31:23

.

12/14/2012 16:36:00Heating with 48 Watts on blue channel. Manual adjustments.Malachi Heder12/14/2012 22:36:00

.

12/14/2012 20:31:15Lowered hydrogen pressure to 7 bar.Malachi Heder12/15/2012 2:31:15

.

12/15/2012 15:36:50Reducing pressure to 6 bar. Pressure seems to be dropping very slowly, still could not find a leak in the cell with the combustible gas detector.Malachi Heder12/15/2012 21:36:50

.

12/16/2012 11:53:16reduced hydrogen pressure to 5 bar.Malachi Heder12/16/2012 17:53:16

.

12/16/2012 17:28:52Reduced hydrogen pressure to 4 bar.Malachi Heder12/16/2012 23:28:52

.

12/16/2012 20:10:36Reducing pressure to 3 bar.Malachi Heder12/17/2012 2:10:36

.

12/16/2012 21:34:55Reduced pressure to 2 bar.Malachi Heder12/17/2012 3:34:55

.

12/16/2012 23:00:10Reduced pressure to 1 bar.Malachi Heder12/17/2012 5:00:10

.

12/17/2012 9:27:05Cooling off cell.Malachi Heder12/17/2012 15:27:05

.

12/17/2012 9:54:13Increasing pressure to 6.3 bar to get 8 bar when hot.Malachi Heder12/17/2012 15:54:132012-12-17

.

12/17/2012 9:58:41Increasing power to 48 Watts.Malachi Heder12/17/2012 15:58:41

.

12/17/2012 12:14:00Reduced Pressure to 7 bar.Malachi Heder12/17/2012 18:14:00

.

12/17/2012 13:22:23Reducing Hydrogen pressure to 6 bar.Malachi Heder12/17/2012 19:22:23

.

12/17/2012 14:39:41Reduced pressure to 5 bar.Malachi Heder12/17/2012 20:39:41

.

12/17/2012 17:04:28Reduced pressure to 4 bar.Malachi Heder12/17/2012 23:04:28

.

12/17/2012 19:56:53Reduced pressure to 3 bar.Malachi Heder12/18/2012 1:56:53

.

12/17/2012 20:34:04Pressure reduced to 2 bar.Malachi Heder12/18/2012 2:34:04

.

12/17/2012 22:14:47Reduced to 1 bar.Malachi Heder12/18/2012 4:14:47

.

12/18/2012 9:05:31Turned cell off, cooling before power range sweep.Malachi Heder12/18/2012 15:05:31

.

12/18/2012 10:04:11Turned on Blue Channel. Script running that will step up power from min to max and back down to see if the temperatures equilibrate in 45 min. We will compare the increasing equilibrium to the decreasing equilibrium at each step.Malachi Heder12/18/2012 16:04:11

.

12/18/2012 10:35:07Turning script off, we are working on the T_GlassOut sensor.Malachi Heder12/18/2012 16:35:07

.

12/18/2012 11:16:47Changed out T_GlassOut sensor by adding aluminum and thermal grease to the outside of the glass. Now starting the script over again.Malachi Heder12/18/2012 17:16:47

.

12/19/2012 10:04:204:00 pm 12/18/2012Added 8 new thermocouples to Celani Cell #1 and surrounding area. With the addition of new thermocouples we created a new test. Info regarding placement and naming scheme is on the way. Working on getting the test to update properly currently.Malachi Heder12/19/2012 16:04:20

.

12/19/2012 11:49:44The new test for temperatures on and around the cell is updating to the database. It will take a few hours to update to current readings.Malachi Heder12/19/2012 17:49:44

.

12/19/2012 12:07:32To avoid much confusion the new 8 sensors test has been renamed to: Celani Cell #1: Extra Cell and Area Temperatures. Malachi Heder12/19/2012 18:07:32

.

12/19/2012 14:48:18We are adjusting the database for the new test with the extra thermocouples. We will have it figured out shortly.Malachi Heder12/19/2012 20:48:18

.

12/20/2012 9:44:30Testing T_A3 in boiling water to determine accuracy. It is a new type of thermocouple.Malachi Heder12/20/2012 15:44:30

.

12/20/2012 10:20:59Putting all three Ambient sensors (T_A1-3) into a copper pipe, isolated from each other to see how precise the three are.Malachi Heder12/20/2012 16:20:59

.

12/20/2012 11:42:48We are disassembling the cell for vertical mounting with the intent of running calibrations for the whole holiday weekend.Ryan Hunt12/20/2012 17:42:48

.

12/21/2012 13:07:48Just bringing Cell 1.0 and 1.1 online. Blog entry to follow shortly to explain the new set up
Ryan Hunt12/21/2012 19:07:48

.

12/21/2012 13:36:25Vacuuming down cells to test tightness of everything. Achieved 160 mTorr with both cells vacuuming simultaneously.Ryan Hunt12/21/2012 19:36:25

.

12/21/2012 15:04:19Apparatus checks out. The data collector is operating. The cells are tight to 100 milliTorr and slowly dropping as it nears the limits of this vacuum pump. Time to load with 0.5 bar H2 and run a calibration cycle or two between now and tomorrow evening. The plan is to up to 1 bar and then run that cycle repeatedly till Dec 26th.Ryan Hunt12/21/2012 21:04:19

.

12/21/2012 15:13:37Cells loaded to 0.5 Bar. They were connected to eachother at the time of hte loading, but we see .4765 bar in one and .5010 bar in the other. Variation between pressure sensors, I guess. Ryan Hunt12/21/2012 21:13:37

.

12/21/2012 15:26:49Stepping calibration scripts started on both cells. The blue channel is the NiChrome wire. The Red channel is the bare isotan. The calibration is running on the Isotan wires.
The calibration steps up to about 40 volts in 7 steps and then back down to zero. Each cycle should take 11 hours and 15 minutes.
Ryan Hunt12/21/2012 21:26:49

.

12/21/2012 15:56:48The T_Air 1, 2 and 3 sensors are not hooked up at the moment. Will try to get that almost immediatelyRyan Hunt12/21/2012 21:56:48

.

12/21/2012 21:32:50Fixed T_Ambient on the Cell 1.0 test. It was pointed to the wrong sensor input. We will be able to use the T_ambient from the other test to back fill the data for the first part of the test. They both are referencing the same sensor.Ryan Hunt12/22/2012 3:32:50

.

12/21/2012 21:56:21Finished hooking up T_A1, T_A2, and T_A3 again. I had to rearrange which was wired to what inputs and them made the appropriate changes in the Test definition.Ryan Hunt12/22/2012 3:56:21

.

12/22/2012 15:12:36After completing 2 full calibration cycles starting at 0.5 bar H2, Just upped the pressure to 1 bar of H2. The cells will run this way for 6 or 7 cycles (through Dec 25th) We will look at the variation over that time and see how consistent it is run from run.Ryan Hunt12/22/2012 21:12:36

.

12/26/2012 15:09:4312/26/2012 9:00:00I turned off the calibration script to let the test cells sit at high temperature for a few hours and see how the equilibrium temperatures looked.Ryan Hunt12/26/2012 21:09:43

.

12/26/2012 15:22:32Just Shut off Cell 1.0 to let it cool. I want to see what effect this has on the cell temps in cell 1.1 next door. Ryan Hunt12/26/2012 21:22:32

.

12/26/2012 17:07:50Just turned on power to cell 1.0 again. The data from the cool down was lost on the data collector because it had locked up. We still have a before and after point to compare.Ryan Hunt12/26/2012 23:07:50

.

12/27/2012 9:01:56Flipped off cell 1.1 to let it cool. will rewarm in a while do we can test time constantsRyan Hunt12/27/2012 15:01:56

.

12/27/2012 9:49:029:00:00T_GlassOut sensor "popped" out from behind the aluminum/grease plate. I bent the thermocouple so that there is no spring tension pushing it out. I inspected the T_GlassIn thermocouple for good contact. It looks unchanged and still in good contact like it was at the beginning of the calibration runs.Malachi Heder12/27/2012 15:49:02

.

12/27/2012 11:37:23Turned Cell 1.1 back on after a full cooling.Ryan Hunt12/27/2012 17:37:23

.

12/27/2012 18:19:43Vacuuming/adding H2 to 0.8-0.9 bar when hot in cells 1.0 and 1.1.Malachi Heder12/28/2012 0:19:43

.

12/27/2012 18:22:51Starting 0.5 bar H2 calibration run in both cells.Malachi Heder12/28/2012 0:22:51

.

12/28/2012 15:55:02Loading both cells to just over one bar of H2 and testing for leaks.Malachi Heder12/28/2012 21:55:02

.

12/28/2012 16:10:21No detectable leak in version 1.1. Leak in pass though in version 1.0 (same as before).Malachi Heder12/28/2012 22:10:21

.

12/28/2012 16:37:57Vacuuming both cells and cooking out hydrogen. Gearing up for a helium run.Malachi Heder12/28/2012 22:37:57

.

12/28/2012 17:01:52In our haste to prep the cells for helium calibrations, we forgot to lower the power on each wire while we vacuumed them. They became red hot and coated the inside of the quartz/borosilicate. They burnt.Malachi Heder12/28/2012 23:01:52

.

12/29/2012 8:47:41Stopping tests for cleaning of reactors.Malachi Heder12/29/2012 14:47:41

.

12/30/2012 11:31:02Apparatus has been rebuilt and is back together. We are recommissioning it and making sure all the sensors work and that it holds a vacuum.Ryan Hunt12/30/2012 17:31:02

.

12/30/2012 13:07:54Helium added to both cells. T_Mica ~ 50 C, waiting 0.5 hours until T_Mica drops to room temperature. Ryan Hunt12/30/2012 19:07:54

.

12/30/2012 14:26:41Started 1.0 bar helium calibration cycles on both cells for comparison with Celani Wire after it is installed. Ryan Hunt12/30/2012 20:26:41

.

12/31/2012 11:48:39Cell 1.0 turned off and decommissioned because of black fuzzy accumulation on the Isotan 44 wire.Malachi Heder12/31/2012 17:48:39

.

12/31/2012 16:06:16Replaced Isotan wire, reassembled and installed the cell. Tested for Vacuum tightness. Loaded with 1.0 bar He. Started up and down step calibration cycles, again.Ryan Hunt12/31/2012 22:06:16

.

1/2/2013 9:23:32Terminated the script and letting the temperatures hold in order to cook out stuff. Also noted a light hazy film at the very bottom of cell 1.1, which has the Macor ceramic. No idea what that may be,Ryan Hunt1/2/2013 15:23:32

.

1/2/2013 11:09:56Pressure is rising almost .15 bar in 2 hours as more contaminant material breaks down into gas and black residue on the wire. Ryan Hunt1/2/2013 17:09:56

.

1/3/2013 9:00:41Reduced the pressure in cell 1.0 in order to see if any more off-gassing occurred at lower pressures. Only went down to .98 bar.Ryan Hunt1/3/2013 15:00:41

.

1/3/2013 9:08:08Barometric pressure right now is 30.21 inches of mercury = 1.0229106 bar
-Just in case we suspect leaking into the cell
Ryan Hunt1/3/2013 15:08:08

.

1/3/2013 13:04:00Turning off both cells to open and clean. Cell 1.0 needs another new isotan wire to replace the gunked up one. Cell 1.1 has a tiny patch of hazy condensate on the glass at the bottom of hte cell.Ryan Hunt1/3/2013 19:04:00

.

1/3/2013 16:15:45Cells cleaned, new Isotan in Cell 1.0. Testing for Vacuum tightness.Ryan Hunt1/3/2013 22:15:45

.

1/3/2013 16:52:52Loaded both cells with 1.0 bar He and started the same calibration script, again. The intent ot make sure we don see anymore fouling and to add to our statistical data about runs in He before we do a Hydrogen run to compare against the ones we did before.
Ryan Hunt1/3/2013 22:52:52

.

1/4/2013 10:58:25Stopped the helium runs to let it cool and load with hydrogen for some comparative hydrogen runs.Ryan Hunt1/4/2013 16:58:25

.

1/4/2013 13:25:33Both cells turned off and tests stopped. Troubleshooting power supply issues.Malachi Heder1/4/2013 19:25:33

.

1/4/2013 14:06:41Problem with power supplies on both the test cells. The software tured the power off too quickly and damaged the FET on the control interface board that pulled the power to zero. The workaround was to cut the FET out. The tradeoff is the power supply won't go to zero volts when told to. It will bottom out at about 1 volt.
After we got that fixed, we started the calibration process for 1.0 bar H2 at 1:55 pm.
Ryan Hunt1/4/2013 20:06:41

.

1/8/2013 9:40:491/7/2013 9:00:00The data collector locked up from 9am to 7 pm, yesterday, or so. We lost data in that time frame. It was not important, thoughRyan Hunt1/8/2013 15:40:49

.

1/8/2013 10:28:52Turning off power to both cells. Once cooled we will be putting in Celani wire.Malachi Heder1/8/2013 16:28:52

.

1/8/2013 10:44:28Updating software while cells are cooling down.Malachi Heder1/8/2013 16:44:28

.

1/8/2013 11:31:04Removing vertical cells to put Celani wire in them! :)Ryan Hunt1/8/2013 17:31:04

.

1/8/2013 15:28:20Cells are hooked up and tests are started. We will be vacuuming both cells for a few hours.

Cell 1.0: 0.2g of 14 layer Celani wire, 95 cm long, measured out of bag at 16.5 Ohms.

Cell 1.1: 0.2g of 2 layer Celani wire, ~95 cm long, measured out of bag at 16 Ohms.
Malachi Heder1/8/2013 21:28:20

.

1/8/2013 16:33:33Cell 1.0 is not as vacuum tight as we would have hoped. Taking it apart to replace the copper gasket.Malachi Heder1/8/2013 22:33:33

.

1/8/2013 17:13:48Hooking up Cell 1.0 to test vacuum again.Malachi Heder1/8/2013 23:13:48

.

1/9/2013 8:51:20Turning off vacuum. It achieved somewhere near 10⁻5 torr after setting all night. The vacuum line is long and narrow, so it pumps down very slow. Those are good tight cells.Ryan Hunt1/9/2013 14:51:20

.

1/9/2013 9:12:25Starting a short run in helium. Four steps up and four steps down with Celani wire.Malachi Heder1/9/2013 15:12:25

.

1/9/2013 9:37:55Changing sensor layout in the tests for Cell 1.0 and 1.1Malachi Heder1/9/2013 15:37:55

.

1/9/2013 9:41:26I have changed T_Mica to T_Macor in cell 1.1 and T_Rise to T_Macor Rise. Also, disregard P_Out and P_Xs for now. We are working on these sensors currently.Malachi Heder1/9/2013 15:41:26

.

1/9/2013 9:58:04I have added new fields to cell 1.0 to look more like cell 1.1 (P_In, P_Out, P_Xs, T_Mica Rise). Again we are working on P_Out and P_Xs.Malachi Heder1/9/2013 15:58:04

.

1/9/2013 10:27:28The correct values are in place for the calculated P_Out and P_Xs in both tests.Malachi Heder1/9/2013 16:27:28

.

1/9/2013 18:14:32Powering up to the highest step to check for consistent temperature readings.Malachi Heder1/10/2013 0:14:32

.

1/9/2013 22:29:19Troubleshooting instrumentation to try to rule out instrumentation error because the cell 1.0 was reading 10c too low. Any reading may be affected by our troubleshooting tlll we give the all clear.Ryan Hunt1/10/2013 4:29:19

.

1/9/2013 23:34:03Turned on the same calibration steps as earlier this morning after renewing helium pressure inside each cell.Malachi Heder1/10/2013 5:34:03

.

1/10/2013 15:10:56Updating firmware and software, downtime expected along with wrong values for short periods.Malachi Heder1/10/2013 21:10:56

.

1/10/2013 15:39:35Powering up cells to check sensors with new firmware.Malachi Heder1/10/2013 21:39:35

.

1/11/2013 9:46:02Bad connection on T_GlassIn in cell 1.0. It has been removed from the test.Malachi Heder1/11/2013 15:46:02

.

1/11/2013 9:48:06Switching both cells to heating with the passive (NiChrome) wire to study radiation profile on the cells. Thermal images/movie will follow.Malachi Heder1/11/2013 15:48:06

.

1/11/2013 10:43:33Cell 1.0 test is down while the database catches up to current readings.Malachi Heder1/11/2013 16:43:33

.

1/11/2013 11:34:59Switching cell 1.0 from NiCr to Celani wire at the same power level. David Roberson, this is for you.Malachi Heder1/11/2013 17:34:59

.

1/11/2013 12:52:16Shutting down cells to cool before beginning hydrogen loading phase. Cell 1.0 may just be an invalid experiment, but we will try to load and see how it behaves in hydrogen. We are frustrated with this experiment design and there is no proof we could get this apparatus to work better if we delayed and fiddled with it. Onward to the experiment and then to new experiment designs.Ryan Hunt1/11/2013 18:52:16

.

1/11/2013 14:06:59Vacuuming both cells.Malachi Heder1/11/2013 20:06:59

.

1/11/2013 14:30:16Loading with Hydrogen, approx 1 bar.Malachi Heder1/11/2013 20:30:16

.

1/11/2013 15:10:48Kicked on the power to 50 W approx. to begin the loading phase on both cells. Yippee. Here we go!Ryan Hunt1/11/2013 21:10:48

.

1/11/2013 16:23:51Upping the power to get a higher temp and try to get cell 1.1 to drop resistance again after leveling off. Going to up both cells, though.
Ryan Hunt1/11/2013 22:23:51

.

1/12/2013 11:27:04Loading at this temp seems to have leveled off. We will sweep from 200 C to roughly 300 C and see if we observe more loading at different temps. Still heating with the nicr wire. 14 30 minute steps for about 7 hours of more loadingRyan Hunt1/12/2013 17:27:04

.

1/12/2013 20:05:45Letting the cells cool down after loading. Ryan Hunt1/13/2013 2:05:45

.

1/12/2013 23:04:56Going live - shooting for 1.8 amps through each wire
Cell 1.0 Estimated impedance of 15.2 ohms at temp
1.8 Amps X 15.2=27.36V=49.25 W
DAC setting of Cell 1.1 Estimated resistance of 16.3 ohms at temp
1.8 Amps X 16.3=29.34V=52.8 W
But we'll step into it in a few steps
Ryan Hunt1/13/2013 5:04:56

.

1/13/2013 13:09:19Just troubleshot a bit of noise in the voltage signal on cell 1.1. Looks like it was either a loose connection or an inconsistent contact in the switch on the power supply unit. It appears that the voltage rose after that back to where it had started.

I am now going to turn on the blue lines (the Nichrome wires) to get each cell hotter. If that doesnt do it, we will add an independent power supply for the blue line so we can control it independently. Right now there is only one power supply per cell.
Ryan Hunt1/13/2013 19:09:19

.

1/13/2013 18:32:3618:20 to 18:22 US Central time - voltage drop on red, but not blue. bad connectin issue again. caused what looked like a small P_xs rise. Time to check some connections.Ryan Hunt1/14/2013 0:32:36

.

1/13/2013 18:38:33Pretty sure I got it. Loose connection where the wire going into the cell connected to the first pin of the terminal strip on the connector plate. The wire fit just to the side of the screw that is supposed to pin it. Seems tight, now.Ryan Hunt1/14/2013 0:38:33

.

1/14/2013 10:25:19Shutting down cells to install independent power supplies for the Nicrhome wire on the blue channel for both cells.Ryan Hunt1/14/2013 16:25:19

.

1/14/2013 11:36:20Second power supplies installed and seem to be working. They are manual adjust only right now.Ryan Hunt1/14/2013 17:36:20

.

1/14/2013 13:11:22Set the Heating wires to 60W. Now we can heat up the Celani wire and see what we see.
Ryan Hunt1/14/2013 19:11:22

.

1/14/2013 13:20:00Gonna step the power up gradually in about 30 5 minute steps up to right near the 1.8 amp limit. Then adjust from there. Hoping weĺl see a reaction start at a particular limit.Ryan Hunt1/14/2013 19:20:00

.

1/15/2013 16:09:03Just upped the voltage a tiny bit on each one. Ryan Hunt1/15/2013 22:09:03

.

1/16/2013 9:29:12sweeping both cells up ~0.05Amps in 10, 10 minute steps.Malachi Heder1/16/2013 15:29:12

.

1/17/2013 10:04:35Continuing stepping down in current ~0.04 amps. 45 min steps.Malachi Heder1/17/2013 16:04:35

.

1/17/2013 13:42:12Going back to the areas from last night where the excess power was highest. Oscillating in 30 min steps up and down through that range. Cell 1.0 and Cell 1.1 have different ranges.Malachi Heder1/17/2013 19:42:12

.

1/18/2013 14:02:29We raised the pressure in each cell to be closer to the calibration curves.

Cell 1.0: shooting for 1.54 bar
Cell 1.1: shooting for 1.38 bar

Will watch as it drops through the calibration curve ranges.
Malachi Heder1/18/2013 20:02:29

.

1/20/2013 8:27:27Made a visual inspection of the apparatus. The wires appear pristine. The Celani wires are darker grey than the nichrome and maybe have a slight copper color. There appeared to be no shorts in the wire. I had suspected the wires shifted slightly and contacted the copper terminal wire near the end as a cause for the slight but sudden resistance drop on the blue channel in cell 1.1. I could not see any evidence of that. The rest of the system looked no different, either. Ryan Hunt1/20/2013 14:27:27

.

1/21/2013 13:59:34At 11:42 AM the door to the lab was left open for approximately 30 seconds. This caused a downward spike in the ambient temperature.Malachi Heder1/21/2013 19:59:34

.

1/22/2013 9:34:27Computer was down from 8:00 am to 9:33 am. Working to get it back up and running.Malachi Heder1/22/2013 15:34:27

.

1/22/2013 9:46:31The computer is back online. Looks like a potential hard drive problem.Malachi Heder1/22/2013 15:46:31

.

1/22/2013 10:17:36Letting the cells cool down. A hard drive error has plagued our machine. We will resume after a bit of trouble shooting and see if everything comes back to the same level.Malachi Heder1/22/2013 16:17:36

.

1/22/2013 16:48:36We are currently testing the integrity of our hard drive. We will resume testing tomorrow morning after we have the steps determined for David Roberson's calibration/run.

Steps are summarized as follows (settling time will be 1.5 hours and we will do our best to step abruptly with as little fine tuning as possible):

1. Celani wire at zero power
2. Celani wire at TBD power
3. Celani wire at TBD power
4. Celani wire at Max power

5. Heating wire at TBD power with Celani wire at Max power
6. Heating wire at TBD power with Celani wire at Max power
7. Heating wire at TBD power with Celani wire at Max power
8. Heating wire at Max power with Celani wire at Max power

9. Cooling off to room temp

10. Step from zero to max power on Celani wire
11. Step from zero to max power on Heating wire while Celani wire is at Max power

This is a tentative plan and is subject to change between now and tomorrow morning.
Malachi Heder1/22/2013 22:48:36

.

1/22/2013 17:00:08Tests are back up and running. No errors in our hard drive test. Again, we will resume active runs tomorrow morning.Malachi Heder1/22/2013 23:00:08

.

1/23/2013 8:39:02Vacuuming both cells.Malachi Heder1/23/2013 14:39:02

.

1/23/2013 10:31:54Starting the calibration run. Steps on the celani wire (Red channel) are:

0W
15W
30W
50W
Malachi Heder1/23/2013 16:31:54

.

1/23/2013 11:13:03Test was started at ~10:30 am local time.Malachi Heder1/23/2013 17:13:03

.

1/23/2013 13:44:51The dip in T_ambient was real. The door was blown open over our lunch period and was closed after.Malachi Heder1/23/2013 19:44:51

.

1/24/2013 9:07:11Set the blue channels to approx 12 W on each cell for the next step in this test. The blue channels are manually adjusted and not scriptable, yet, so we couldnt do it overnight.Ryan Hunt

.

1/24/2013 10:36:40Second step manually adjusted
Ryan Hunt

.

1/24/2013 12:12:56Third step done.Ryan Hunt

.

1/24/2013 16:37:03Cells are turned off and letting them cool before the last two big steps.Malachi Heder

.

1/24/2013 16:41:52Red channels will come on in 1.5 hours and then in 3 hours I will turn the blue channels on.Malachi Heder

.

1/24/2013 19:48:44final step with the blue channel.Malachi Heder

.

1/25/2013 10:37:28Stepping up by .1 amps from 1.8 to 1.85 on celani wire, 45 min steps. Changed NiCr (blue heating wire) to 55 watts.Malachi Heder

.

1/28/2013 9:02:25Raising the current in each cell to 1.86 amps. Cell 1.0 will have 6 steps and cell 1.1 will have 4 steps.Malachi Heder

.

1/28/2013 13:59:24We've been tracking data errors that show zero on some fields. The problem was caused by one acquisition board that had a bad data connection. That caused the polling script to lockup and cause up to 40 seconds between reads, or a failure. The results are NULLs in the database, but wrongly displayed as zeros on the data table and in the graph. That piece of the bug is still being fixed, and will allow the graph to look a lot better.Ryan Hunt

.

1/28/2013 14:05:2813:45:00For cell 1.1, the blue power supply was bumped slightly causing a small increase in power of under 0.2W. This happened as Malachi was trouble shooting the data acquisition card. That power supply is manually adjusted and a slight bump might have tweaked the potentiometer that controls it.Ryan Hunt

.

1/31/2013 10:52:17The ambient rise at about 10:10 local time is due to a "bake-out" of a few mica pieces. It will continue for a few hours.Malachi Heder

.

2/6/2013 9:28:47Data from this morning and last night has a resolution of 40 seconds, it was due to a service error. Fixed at around 9 local time.Malachi Heder

.

2/6/2013 12:44:48We will be updating the software and firmware. Expect a downtime of 10 minutes or so. During this time the power on the celani wire in both cells will drop to zero for a short time. Hopefully we won lose too much heat.Malachi Heder

.

2/6/2013 13:01:46Updates are done. We lost a few degrees, but everything should come right back up to steady state soon enough.Malachi Heder

.

2/6/2013 13:44:00We are also updating the live data user interface with a few new features.Malachi Heder

.

2/6/2013 14:04:48A couple changes you can enjoy on the Live Data user interface are:

Having multiple test tabs open in the same window and being able to use the toggle button on both of them without interference from the other.

A few more time frames to look at live (updating data) in the drop down menu next to PERIOD LENGTH.
Malachi Heder

.

2/6/2013 15:13:08We see that the pressure in cell 1.0 did not come back up to where it was before we updated the firmware. It doesn't seem to be a software or hardware related issue or, for that matter, a false reading at all.Malachi Heder

.

2/7/2013 10:39:00I took off the aluminum band over the T_GlassOut sensor and cleaned it off. Then I put a small amount of thermal paste on the glass side and put the thermocouple back in the correct position. I did this because the thermocouple was loose underneath the aluminum plate and by touching it very lightly it decreased the temperature by 4 degrees. This fix should stabilize the sensor and create better contact with the aluminum and glass.Malachi Heder

.

2/11/2013 15:15:23While taking pictures in the past 15 minutes, I accidentally loosened a connection to the red channel resulting in a drop in power for a few minutes. It has since been fixed.Malachi Heder

.

2/12/2013 10:25:42Cooling for 5 minutes then heating back up.Malachi Heder

.

2/12/2013 13:51:02Turning off blue channel.Malachi Heder

.

2/13/2013 8:51:17Turing the blue channel back on and letting it settle. Up next is a full cooldown and reheating.Malachi Heder

.

2/13/2013 11:38:19Doing another 5 min cooldown on both cells to see if the loose wire was the cause of the resistance spike in cell 1.1.Malachi Heder

.

2/13/2013 13:41:27Unscheduled down time on the server. Working to fix ASAP.Malachi Heder

.

2/13/2013 14:27:05Server is back up and running, it will take a little while to catch up to the data.Malachi Heder

.

2/13/2013 14:29:25Running another 5 minute cool down, this time from the control software.Malachi Heder

.

2/14/2013 9:42:10Starting the full shutdown now.Malachi Heder

.

2/14/2013 11:09:34Turing the cells back on now.Malachi Heder

.

2/14/2013 13:51:55Raising H2 pressure to 1.54 bar in cell 1.0 and 1.40 bar in cell 1.1. This will be done carefully while the cells are warm.Malachi Heder

.

2/15/2013 9:26:23Turning off the cells, then vacuuming and filling with Hydrogen to cell 1.0 = 1.54 bar and cell 1.1 = 1.54 bar.Malachi Heder

.

2/15/2013 10:47:53I was mistaken. Since the cell will be cooled off, we will raise cell 1.0 to 0.943 bar and cell 1.1 to 0.891 bar. This will be after vacuuming.Malachi Heder

.

2/15/2013 11:28:56We are going to power up the cells in two steps, the first is only the red channel and the second is both the red and blue channels.Malachi Heder

.

2/16/2013 13:56:17Turning on the blue channel to get cell 1.0 to ~105 watts and cell 1.1 to ~103. This will allow direct comparison to calibration runs.Malachi Heder

.

2/18/2013 10:19:19Final adjustments on Blue Channel power.Malachi Heder

.

2/22/2013 9:53:46Pressurizing to 1.5 bar in each cell. Keeping the cells above atmospheric.Malachi Heder

.

2/26/2013 14:05:48Raising the pressure to 1.54 in each cell.Malachi Heder

.

2/28/2013 13:28:43Topping off the reactors with hydrogen. 1.5 bar each.Malachi Heder

.

3/4/2013 6:23:27Loading hydrogen back to 1.5 bar each.Malachi Heder

.

3/5/2013 13:57:46Loading both cells to 1.5 barMalachi Heder

.

3/5/2013 14:12:46H2 sense tower disconnected from supply manifold to move to the other corner of the LENR lab. H2 supply control and power, hood fan control and power, and H2 sense equipment was temporarily disconnected to allow for this change.Wes

.

3/5/2013 16:32:57Done rearranging out in the lab area. Last few hours you will see how the presence of a few human bodies effects the cells.Malachi Heder

.

3/6/2013 9:12:44There will be more activity in the lab, we are making a bench to create more desktop area.Malachi Heder

.

3/11/2013 11:54:18Pressurizing to 1.5 bar each. Malachi Heder

.

3/11/2013 14:32:27In the middle of the week we will be putting up more counter space. We will be moving the dual vertical cells out from under the hood and on to the new counter space. This will require down time and the cells will cool off. I will let you know when we start working in there.Malachi Heder

.

3/12/2013 15:10:51The cells will be shut down, moved and we will upgrade the software. This will take a few days. This is preparation for the new cells that will be commissioned soon.Malachi Heder

.

3/13/2013 15:36:18The cells are unhooked while we work on them.Malachi Heder

.

3/18/2013 15:51:59Powering up Cell 1.0 and 1.1.Malachi Heder

.

3/19/2013 8:57:35I turned on the fan in the lab to stir up the air around the cells, it is not blowing directly on them. Since we moved these cells out of the hood, the air currents are different that when we had calibrated the cells.Malachi Heder

.

3/21/2013 8:16:24Had to restart the power in the Celani Wire (Red). It cooled off a little.Malachi Heder

.

3/22/2013 8:17:30Sorry for the weird data collection going on right now. Our software guy (Scott) is taking a look at it and we should have some idea of why it is happening. This is the same bug from yesterday :(Malachi Heder

.

3/22/2013 9:04:20Sorry again for the delay in fixing this bug. We wanted to figure out exactly what was causing it. It turns out that we had a problem with enabling multiple serial ports for the data collection.Malachi Heder

.

3/22/2013 9:19:34As of now, they are heating up again. Both cells.Malachi Heder

.

3/26/2013 14:08:43Filling cell 1.0 and 1.1 to 1.5 bar H2.Malachi Heder

.

3/27/2013 15:38:41The active cell pulled a vacuum to 3.53E⁻2 mbar. This is the right cell under our hood.Malachi Heder

.

4/17/2013 13:57:44The cells are being shutdown, we have no confidence in the excess energy being displayed on the tests because of the new location and below atmospheric pressure. They will be cooled down and the data flow will be stopped until further notice. A better explanation will be sent out in the blog.Malachi Heder

.

5/16/2013 10:11:29We are recommissioning Cell 1.1 - Macor. We are going to test the Aluminum foil around the glass tube. Pics and detail will be posted on the blog. Malachi Heder

.

7/25/2013 17:49:39Storm took out the power in the lab. Bringing all the experiments back online.Malachi Heder

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25

.

0.25